Consider an economy having a Cobb Douglas production function, where the share of capital income in total income is 1/2. The depreciation rate is , population growth rate is n = 0.02

A. The golden rule level of capital per worker is .

B. The golden rule level of investment per worker is .

C. The golden rule level of output per worker is .

D. The golden rule savings rate is X% where X equals .QUESTION 2 20 poin Consider an economy having a Cobb Douglas production function, where the share of capital income in total
Answer #1

As p.4oduction functim y = 110.5 $0cm sumpling (1--Savírny yale y productim funtim = C (1-5) xy C = y-sy (n+dJK cak 2016 -(0.